A chemist is mixing a solution that is 2% acid and another solution that is 8% acid. He needs to make 75 mL of a solution that is 5% acid. How much of each solution should he use? Solve and write a system of equations to model the situation.
Answer by mananth(16949)

percent percentage ---------------- quantity
Acid type I 2 ---------------- x ml
Acid type II 8 ---------------- 75 - x ml
Mixture 5.00% ---------------- 75

2x+8(75-x)=5*75

2x+600-8x= 375
2x-8x=375-600
-6x=-225
/-6
x = 37.5 ml 2.00% Acid type I
37.5 ml 8.00% Acid type II